The average number of motorcycles which stop at a petrol station is 5.2 per hour. By assuming that the number of motorcycles that stop at the petrol station follows a Poisson distribution, find the probability that


(a) 10 motorcycles stop at the petrol station at an interval of 60 minutes.

(b) more than 3 motorcycles stop at the petrol station at an interval of 10 minutes.


1
Expert's answer
2021-09-09T18:03:59-0400

Mean λ=5.2

"P(X=k) = \frac{λ^k \times e^{-λ}}{k!}"

(a)

"P(X=10)= \frac{5.2^{10} \times e^{-5.2}}{10!} = 0.0219"

(b)

"Mean = \frac{5.2}{6}= 0.866 \\

P(X>3) = 1-P(X≤3) \\

= 1 -[P(X=0) +P(X=1) +P(X=2) +P(X=3)] \\

=1-[ (\frac{0.866^0 \times e^{-0.866}}{0!}) + (\frac{0.866^1 \times e^{-0.866}}{1!}) +(\frac{0.866^2 \times e^{-0.866}}{2!}) + (\frac{0.866^3 \times e^{-0.866}}{3!}) ] \\

= 1 -[e^{-0.866}(1+0.866+0.3749 + 0.1082)] \\

= 1-[0.4206 \times 2.3491] \\

= 1 -0.9880 \\

= 0.0012 \\

P(X>3) = 0.0012"